# Machine Olfaction and Embedded AI Are Shaping the New Global Sensing Industry

### v1

@@ 1 Introduction
Throughout evolution, life has depended on exchanging chemical messages through the air and water. Olfaction, the sense of smell, enables organisms to detect and act upon minute traces of volatile compounds, guiding them toward food and mates, and even giving them the ability to detect human disease by sniffing [DeWeerdt, 2024]. While machine vision and audition now rival the human eye and ear, our technologies have long struggled to match the nose [Covington et al., 2021]. At its core, biological olfaction is a feat of molecular matchmaking and pattern recognition: small molecules called odorants (typically under 350 atomic weight) bind with high specificity to much larger membrane-bound receptors (typically  30,000 molecular weight), triggering even whole-organism behavioral responses to vanishingly small chemical cues. Traditional analytical systems such as Gas Chromatography - Mass Spectrometry (GC-MS), Liquid Chromatography-Mass Spectrometry(LC-MS)LC-MS, and odor fingerprinting approaches involving a variety of sensors have often been attempted as noses. These have traditionally used metal oxide sensors, and interferometric, surface-plasmon resonance (SPR), quartz crystal microbalance (QCM) or THz-time domain spectroscopy (THz-TDS) platforms [Laplatine et al., 2022, Songkhla, 2025, Chapalo et al., 2024, Qin et al., 2022, Koch et al., 2023] and have long achieved remarkable sensitivities under laboratory conditions, but none have led to a machine olfactor with not the contextual nuance, specificity, generalizeability, or speed required to truly mimic the capabilities of the biological sense of smell and the nervous systems that use it.
Two key barriers stood in the way of machine olfaction as it was poised to join the club of machine vision and audition: 1) the “limit of detection” as machine olfactors could not reliably respond to single-molecule events and 2) the “limit of recognition”, the ability to decode complex scent characters in plumes mixed in dynamically fluctuating background noise. Both are being overcome. A breakthrough came with the successful integration of stabilized mammalian olfactory receptors into photonic and electronic devices [Kim et al., 2009, Goldsmith et al., 2011, Yoo et al., 2021a, Kim et al., 2023]. These systems, when married with modern AI, including deep neural networks and large language models [K and Sundararajan, 2025, Boscher et al., 2024, Feng et al., 2025, Liang et al., 2022a] are showing promise in real-world scent recognition.
@@ 2.2 Multisensory Artificial Intelligence
Multisensory artificial intelligence is a vibrant multidisciplinary research field that designs computer agents that can perceive, reason, and interact through multiple communicative modalities, including linguistic, acoustic, visual, tactile, sensory, and physiological messages [Liang et al., 2022a, 2023]. Multimodal AI systems can be of great impact in diverse scientific areas bringing practical benefits, such as supporting human health and well-being [Acosta et al., 2022, Dai et al., 2025a], enabling multimedia content processing [Liang et al., 2021, 2024], and enhancing real-world autonomous agents [Durante et al., 2024, Xie et al., 2024]. Multisensory models trained on combined vision, audio, and olfactory data have already demonstrated superior context interpretation in many medical sensing fields [Dai et al., 2025b, Liang et al., 2022a]. This fusion aligns with a broader AI trend toward embodied cognition and contextual awareness [Duan et al., 2022]. By leveraging advanced multimodal AI techniques, there is potential to combine olfaction with other human senses to improve resilience and inference capability in real-world applications. One example of such a technique is called cross-modal translation [Liang et al., 2022a, Pham et al., 2019], where multimodal models can take in one modality and generate another, such as using a visual image to predict how the substances (whether objects such as food items or odorant molecule identity and concentration descriptors) in the image would feel, taste and/or smell to a human [Brewster et al., 2006, Rombach et al., 2022]. These cross-modal translation models could be trained on paired modalities (e.g., image and its corresponding smell descriptors) to decode chemical or sensor representations of smell. Another technique is called cross-modal fusion [Liang et al., 2022a], which investigates whether jointly learning from images, audio, and scent can improve performance on individual modalities such as visual object recognition, speech recognition, and smell recognition [Huh et al., 2024, Radford et al., 2021, Tjandrasuwita et al., 2025]. Prior work has shown that jointly learning from human verbal and nonverbal communication can improve performance as compared to learning from only verbal spoken words [Ong et al., 2025, Zadeh et al., 2020]. Similarly, today’s large language models are increasingly trained with visual and auditory inputs [Li et al., 2024]. Studying how olfaction can be learned jointly with visual and language modalities can lead to learning algorithms that learn like humans do, grounded in multisensory inputs [Bisk et al., 2020, Liang, 2022]. In addition to multimodal AI models, a key driver for this line of research is multisensory scent environments, where data associate with the same physical space and processes ongoing are collected from multiple sensing modalities [Feng et al., 2025]. For example, collecting unified datasets of foods and beverages with smell sensor readings, photos, text descriptions, chemical molecular analysis etc. merged into generalized “Synesthetic Memory Objects”. These rich multimodal datasets are essential for large-scale multimodal models [Antol et al., 2015, Li et al., 2024, Liang et al., 2022b], and have contributed significantly to today’s multimodal foundation models.

### v2

@@ 1 Introduction
Throughout evolution, life has depended on exchanging chemical messages through the air and water. Olfaction, the sense of smell, enables organisms to detect and act upon minute traces of volatile compounds, guiding them toward food and mates, and even giving them the ability to detect human disease by sniffing [DeWeerdt, 2024]. While machine vision and audition now rival the human eye and ear, our technologies have long struggled to match the nose [Covington et al., 2021]. At its core, biological olfaction is a feat of molecular matchmaking and pattern recognition: small molecules called odorants (typically under 350 atomic weight) bind with high specificity to much larger membrane-bound receptors (typically  30,000 molecular weight), triggering even whole-organism behavioral responses to vanishingly small chemical cues. Traditional analytical systems such as Gas Chromatography - Mass Spectrometry (GC-MS), Liquid Chromatography-Mass Spectrometry(LC-MS), and odor fingerprinting approaches involving a variety of sensors have often been attempted as noses. These have traditionally used metal oxide sensors, and interferometric, surface-plasmon resonance (SPR), quartz crystal microbalance (QCM) or THz-time domain spectroscopy (THz-TDS) platforms [Laplatine et al., 2022, Songkhla, 2025, Chapalo et al., 2024, Qin et al., 2022, Koch et al., 2023] and have long achieved remarkable sensitivities under laboratory conditions, but none have led to a machine olfactor with not the contextual nuance, specificity, generalizeability, or speed required to truly mimic the capabilities of the biological sense of smell and the nervous systems that use it.
Two key barriers stood in the way of machine olfaction as it was poised to join the club of machine vision and audition: 1) the “limit of detection” as machine olfactors could not reliably respond to single-molecule events and 2) the “limit of recognition”, the ability to decode complex scent characters in plumes mixed in dynamically fluctuating background noise. Both are being overcome. A breakthrough came with the successful integration of stabilized mammalian olfactory receptors into photonic and electronic devices [Kim et al., 2009, Goldsmith et al., 2011, Yoo et al., 2021a, Kim et al., 2023]. These systems, when married with modern AI, including deep neural networks and large language models [K and Sundararajan, 2025, Boscher et al., 2024, Feng et al., 2025, Liang et al., 2024a] are showing promise in real-world scent recognition.
@@ 2.2 Multisensory Artificial Intelligence
Multisensory artificial intelligence is a vibrant multidisciplinary research field that designs computer agents that can perceive, reason, and interact through multiple communicative modalities, including linguistic, acoustic, visual, tactile, sensory, and physiological messages [Liang et al., 2024a, 2023]. Multimodal AI systems can be of great impact in diverse scientific areas bringing practical benefits, such as supporting human health and well-being [Acosta et al., 2022, Dai et al., 2025a], enabling multimedia content processing [Liang et al., 2021, 2024b], and enhancing real-world autonomous agents [Durante et al., 2024, Xie et al., 2024]. Multisensory models trained on combined vision, audio, and olfactory data have already demonstrated superior context interpretation in many medical sensing fields [Dai et al., 2025b, Liang et al., 2024a]. This fusion aligns with a broader AI trend toward embodied cognition and contextual awareness [Duan et al., 2022]. By leveraging advanced multimodal AI techniques, there is potential to combine olfaction with other human senses to improve resilience and inference capability in real-world applications. One example of such a technique is called cross-modal translation [Liang et al., 2024a, Pham et al., 2019], where multimodal models can take in one modality and generate another, such as using a visual image to predict how the substances (whether objects such as food items or odorant molecule identity and concentration descriptors) in the image would feel, taste and/or smell to a human [Brewster et al., 2006, Rombach et al., 2022]. These cross-modal translation models could be trained on paired modalities (e.g., image and its corresponding smell descriptors) to decode chemical or sensor representations of smell. Another technique is called cross-modal fusion [Liang et al., 2024a], which investigates whether jointly learning from images, audio, and scent can improve performance on individual modalities such as visual object recognition, speech recognition, and smell recognition [Huh et al., 2024, Radford et al., 2021, Tjandrasuwita et al., 2025]. Prior work has shown that jointly learning from human verbal and nonverbal communication can improve performance as compared to learning from only verbal spoken words [Ong et al., 2025, Zadeh et al., 2020]. Similarly, today’s large language models are increasingly trained with visual and auditory inputs [Li et al., 2024]. Studying how olfaction can be learned jointly with visual and language modalities can lead to learning algorithms that learn like humans do, grounded in multisensory inputs [Bisk et al., 2020, Liang, 2022]. In addition to multimodal AI models, a key driver for this line of research is multisensory scent environments, where data associate with the same physical space and processes ongoing are collected from multiple sensing modalities [Feng et al., 2025]. For example, collecting unified datasets of foods and beverages with smell sensor readings, photos, text descriptions, chemical molecular analysis etc. merged into generalized “Synesthetic Memory Objects”. These rich multimodal datasets are essential for large-scale multimodal models [Antol et al., 2015, Li et al., 2024, Liang et al., 2022], and have contributed significantly to today’s multimodal foundation models.